how many liters of at 10% solution must be mixed with a 50% acid solution in order to obtain 30 liters of a 30% solution?
Answer by stanbon(75887)

how many liters of at 10% solution must be mixed with a 50% acid solution in order to obtain 30 liters of a 30% solution?
-----------
One variable solution:
Equation:
active + active = active
0.10x + 0.50(30-x) = 0.30*30
---
Multiply thru by 100 to get:
10x + 50*30 - 50x = 30*30
---
-40x = -20*30
x = 15 liters (amt. of 10% solution needed in the mixture)
50-15 = 35 liters (amt. of 50% solution needed in the mixture)
--------
Cheers,
Stan H.